OpenStudy (anonymous):

Write the quadratic equation in vertex form. then identify the vertex. y=3x^2-36+101 ?

OpenStudy (amistre64):

the vertex will have a value of x=36/6 = 6 plug 6 into this equation to get the "y" value for it....

OpenStudy (amistre64):

then we can just shuffle it around to get the other equation by those numbers

OpenStudy (amistre64):

y = (x-6)^2 + B; where B is the value you are going to find :)

OpenStudy (anonymous):

Ohh, ok, thanks for the help, i understand now ^.^